# USTC-ICTS/PCFT-25-44

### v2

@@ 2.2 Euclidean path integral near extremality and zero modes
where * denotes complex conjugation. The first order term vanishes as the classical equations of motion are satisfied. The linearized kinetic operator $\bar{\Delta}[\bar{g}]$ in (2.25) depends on the saddle-point as well as the gauge fixing condition. We will use the harmonic gauge condition in the following discussion with the Lagrange density of the gauge fixing term given by
@@
Once the saddle-point $\bar{g}$ is given, the one-loop partition function (2.25) is determined by the eigenvalues of the linearized kinetic operator $\bar{\Delta}[\bar{g}]$. We will use the saddle-point (2.16), which is the near horizon extremal geometry and satisfies the vacuum Einstein equation, to evaluate the one-loop partition function of the extremal accelerating Kerr black hole. The resultant quantum corrections to the black hole partition function help to resolve the mass gap puzzle discussed in subsection 2.1.
The modes $h_{\mu\nu}$ with vanishing eigenvalues of the operator $\bar{\Delta}[\bar{g}]$, which are called zero modes, play crucial roles in the one-loop approximation of the path integral. As we will see in the next subsection, zero modes contribute leading quantum corrections to the partition function in the low temperature limit. Zero modes are usually associated to the background symmetry that are not removed by the gauge fixing. In the harmonic gauge (2.26), zero modes are necessary to satisfy the transverse and traceless conditions and thus are generated by large diffeomorphisms. In other words, a zero mode can be written as a Lie derivative of the background geometry along a certain vector field $\xi$ which has non-trivial behaviour at the boundary. For the saddle-point (2.16), we take the following ansatz for the vector field
@@
which is a second order differential equation of $f_{y}(y)$ and is solved by the following general solution
@@
with integer valued $n$. These vector fields are large diffeomorphisms since they have non-trivial behaviours at the asymptotic boundary $y\to+\infty$
@@
One way to regularize the one-loop partition function is to deform the extremal background by turning on slightly the temperature of the black hole, which in turn will corrects slightly the eigenvalues of the zero modes. In the next subsection, we will use the near extremal configuration given by (2.16) and (2.17) to calculate the leading corrections to the eigenvalues of the zero modes. We will see that these slightly corrected non-vanishing eigenvalues alter the partition function dramatically with a leading low temperature effect.
@@ 2.3 Quantum corrections to the entropy
Note that $\delta\Lambda_{n}$ are small corrections to the eigenvalues which depend linearly on $T$, for non-zero modes with $\bar{\Lambda}_{n}\neq 0$, the corrections are purely polynomial in $T$
@@
and correspondingly produce $\mathcal{O}(T)$ correction to the black hole mass at low temperature. So zero modes are the essential objects to resolve the mass gap puzzle addressed in subsection 2.1.
@@
where we used the fact that $\bar{\Delta}[\bar{g}]$ is self-adjoint, i.e., $\langle h^{(m)}|\bar{\Delta}[\bar{g}]\delta h^{(n)}\rangle=\langle\bar{\Delta}[\bar{g}]h^{(m)}|\delta h^{(n)}\rangle$ and $h^{(n)}$ are orthonormalized. The integration in (2.45) is highly non-trivial. One way to carry out the integrand is to substitute the full perturbed metric $\bar{g}+\delta g$ into the linearized kinetic operator (2.27) and keep the leading $\mathcal{O}(T)$ term. The variation of the operator itself is intractable but it simplifies dramatically when contracting with the zero modes. Here we print the result we find
@@
$T_{q}$ can be viewed as the critical temperature below which the low temperature quantum corrections become important. From (2.49), it is clear that the critical temperature rises as the acceleration of the black hole increases. However, this is not necessarily indicating that the high temperature quantum corrections are significant for a black hole with large acceleration. The mechanism presented here is working for the first order low temperature perturbation theory. The corrections from the levels that beyond the first order are still unknown at present stage.
@@ 3 Accelerating Reissner-Nordström black hole
$m$, $q$, $\alpha$ are the black hole mass, electric charge and acceleration parameters, respectively. $b$ is the time scaling parameter introduced for achieving consistent black hole thermodynamics Anabalon:2018qfv . $\Phi$ is the electric potential at the horizon, which ensures that the electric gauge field $A$ vanishes on the horizon. The accelerating of the charged black hole is induced by the cosmic string attached on the north ($\theta=0$) and south ($\theta=\pi$) poles. The string tensions are represented by the conical defects at each poles. Following the same argument in section 2, the string tensions at the north $\mu_{+}$ and south $\mu_{-}$ poles can be expressed as
@@ 3.1 Near extremal thermodynamics and geometry
The near extremal thermodynamics fucus on the situation that the two horizon radius $r_{+}$ and $r_{-}$ depart a little bit from their extremal value $r_{0}$. Similar to the rotating case, there are three independent thermodynamic quantities which can be chosen as the temperature $T$, electric charge $Q$ and net string tension $\mu$. In the near extremal expansions, we take the temperature as small parameter and keep the electric charge $Q_{0}=r_{0}$ and the net string tension $0<\mu_{0}=r_{0}\alpha<1$ fixed at their extremal values.
@@
The quantum corrections near extremality are introduced by using the Euclidean path integral with saddle-point chosen as the extremal configuration. To find the near horizon near extremal configuration, we therefore perform the following coordinate transformation
@@
The metric (3.15) characterize the near horizon extremal geometry of the accelerating Reissner-Nordström black hole, which is a direct product of AdS2 and S2 with warping factors. (3.17) is the extremal gauge field, which vanishes explicitly on the horizon $y=1$. When combined with the subleading terms (3.17) and (3.18), we get the near horizon near extremal configurations $d\bar{s}^{2}+\delta(ds^{2})$ and $\bar{A}+\delta(A)$ up to $\mathcal{O}(T)$.
@@ 3.2 Euclidean path integral near extremality and zero modes
which necessitates the joint diagonalization of the operators $\bar{\Delta}[\bar{g},\bar{A}]$, $\bar{P}[\bar{g},\bar{A}]$ and $\bar{O}_{\mathrm{int}}[\bar{g},\bar{A}]$. Inspired by the previous accelerating Kerr black hole case, we will focus on the path integral over zero modes of the saddle-points (3.15) and (3.16), which hopefully produces the leading quantum correction to the partition function at low temperature. Terminologically, the zero modes of the matrix operator in (3.32) can be divided into graviton zero modes and photon zero modes.
@@
with an arbitrary constant $C_{1}$, that is regular on the horizon and vanishes for $n=0$. Note that in deriving the above solution, we do not resort to the form of the function $f_{\phi}(y)$, which is different form the rotating black hole case analyzed in subsection 2.2. This is highly due to the near horizon extremal geometry (3.15) is simply a direct product of AdS2 and S2 with warping factors. The function $f_{\phi}(y)$ can be solved independently from
@@
with integer valued $n$. These vector fields are indeed large diffeomorphisms since they does not die off at the asymptotic boundary $y\to\infty$ and thus can not be gauged away. The existence of the two arbitrary constants $C_{1}$ and $C_{2}$ indicates that there are two linearly independent graviton zero modes.
@@
The $n=0,\pm 1$ modes are excluded because the corresponding vector fields are Killing vectors on the background and do not produce fluctuations. The graviton zero modes which are generated by the $\xi$ with $C_{1}=0$ and $C_{2}\neq 0$ are call vector modes, whose explicit forms can be written as
@@
Note that for the vector modes, we have $|n|\geq 1$. The $n=0$ mode vanishes due to the fact that $\mathop{}\!\partial_{\phi}$ is the rotational Killing vector of the background. One can check that both (3.40) and (3.42) belong to the kernel of the operator $\bar{\Delta}[\bar{g},\bar{A}]$. It is worth to note that in the absence of acceleration, the Reissner-Nordström black hole has three sets of vector modes since there are three independent Killing vectors on S2 Banerjee:2023quv ; Blacker:2025zca . The presence of acceleration deforms the S2 and kills two of the Killing vectors, leaving only the translational symmetry along $\phi$. So there is only one set of the vector modes (3.42) in the accelerating case.
@@
It is straightforward to check that (3.47) is in the kernel of the operator $\bar{P}[\bar{g},\bar{A}]$. For a torsion free theory, the photon zero modes determined by (3.44) always give symmetric $\bar{\nabla}_{\mu}a_{\nu}$. It is obvious that the mixing term (3.31) vanishes for such zero modes. So the previous derived tensor, vector and photon zero modes are indeed belong to the kernel of the matrix operator in (3.32).
@@ 3.3 Quantum corrections to the entropy
This will perturb the operators $\bar{\Delta}$ and $\bar{P}-(\bar{O}_{\mathrm{int}}^{\dagger}\bar{\Delta}^{-1}\bar{O}_{\mathrm{int}})/4$ and correct the the eigenvalues of the zero modes to nonzero ones. According to the perturbation theory at first order, the eigenvalue correction can be evaluated as the expectation value of the perturbed operator with respect to the original orthonormal eigenstates. However, from (3.31), it is obvious that the photon zero modes (3.47) satisfies
@@
So the variation of the operator $(\bar{O}_{\mathrm{int}}^{\dagger}\bar{\Delta}^{-1}\bar{O}_{\mathrm{int}})/4$ will not contribute to the first order eigenvalue correction. Thus, to figure out the eigenvalue corrections, we only need to consider the variations of the operators $\bar{\Delta}$ and $\bar{P}$. Given the orthonormalized zero modes (3.40), (3.42) and (3.47), the corrections to their eigenvalues can be formulated as
@@
$T_{q}^{t}=\frac{r_{0}(3+\mu_{0}^{2})}{(1-\mu_{0}^{2})\sqrt{1+\mu_{0}^{2}}}\,,~~T_{q}^{v}=\frac{r_{0}(1-\mu_{0}^{2})(\mathrm{arctanh}\mu_{0}-\mu_{0})}{\mu_{0}^{3}\sqrt{1+\mu_{0}^{2}}}\,,~~T_{q}^{p}=\frac{r_{0}\mu_{0}}{\left(\mu_{0}-(1-\mu_{0}^{2})\mathrm{arctanh}\mu_{0}\right)\sqrt{1+\mu_{0}^{2}}}\,$, (3.63)
@@
where $S_{0}$ is the extremal entropy. Note that the temperature dependence in the leading term of (3.64) is different from that in the accelerating Kerr case (2.50). This is because the absence of rotation and the presence of the gauge field each leads to additional zero modes and contributes $T^{1/2}$ factor in the partition function. Accordingly, the black hole thermal energy above extremality in this case is dominated by a term scales as $(5/2)T$ at low temperature which replace it on the same footing as the Hawking radiation.
@@ 4 Accelerating Kerr-Newmann black hole
We finally analyze the low temperature quantum corrections to the thermodynamics of the C-metric with both rotation and charge. As we will see, by turning on the rotation, the gravitational zero modes of vector type for the accelerating Reissner-Nordström black hole are no longer the case. The spacetime line element and the $U(1)$ gauge field for the accelerating Kerr-Newmann black hole in the Boyer-Lindquist coordinate system take the following forms Hong:2004dm ; Griffiths:2005se ; Griffiths:2005qp
@@
$m$, $a$, $q$ and $\alpha$ are the black hole mass, angular momentum, charge and acceleration parameters, respectively. $b$ is the time scaling parameter and $\Phi$ is the charge potential at the horizon. The outer and inner horizon radius $r_{+}$ and $r_{-}$ are now given by
@@ 4.1 Near extremal thermodynamics and geometry
As a consistency check, the chargeless limit $q_{0}\to 0$ of the above expansions recover the accelerating Kerr black hole case (2.10), (2.11), (2.12) and (2.13) while the static limit $q_{0}\to r_{0}$ reproduce the accelerating Reissner-Nordström black hole case (3.10), (3.11), (3.12) and (3.13).
To compute the low temperature quantum corrections to the thermodynamic quantities, we need to figure out the near extremal configuration which can be obtained by the following coordinate transformation
@@
When substituting the transformations (4.14), (4.15) and (4.16) into the metric (4.1) and the gauge field (4.2) and taking the small temperature limit with fixed angular momentum, charge and net string tension at their extremal values, we find the leading $\mathcal{O}(T^{0})$ terms

### v3

@@ 2.2 Euclidean path integral near extremality and zero modes
where * denotes complex conjugation. The first-order term vanishes as the classical equations of motion are satisfied. The linearized kinetic operator $\bar{\Delta}[\bar{g}]$ in (2.25) depends on the saddle-point as well as the gauge fixing condition. We will use the harmonic gauge condition in the following discussion with the Lagrange density of the gauge fixing term given by
@@
Once the saddle-point $\bar{g}$ is given, the one-loop partition function (2.25) is determined by the eigenvalues of the linearized kinetic operator $\bar{\Delta}[\bar{g}]$. We will use the saddle-point (2.16), which is the near horizon extremal geometry and satisfies the vacuum Einstein equation, to evaluate the one-loop partition function of the extremal accelerating Kerr black hole. The resultant quantum corrections to the black hole partition function help to resolve the mass gap puzzle discussed in Sec. 2.1.
The modes $h_{\mu\nu}$ with vanishing eigenvalues of the operator $\bar{\Delta}[\bar{g}]$, which are called zero modes, play crucial roles in the one-loop approximation of the path integral. As we will see in the next subsection, zero modes contribute leading quantum corrections to the partition function in the low-temperature limit. Zero modes are usually associated to the background symmetry that are not removed by the gauge fixing. In the harmonic gauge (2.26), zero modes are necessary to satisfy the transverse and traceless conditions and thus are generated by large diffeomorphisms. In other words, a zero mode can be written as a Lie derivative of the background geometry along a certain vector field $\xi$ which has nontrivial behavior at the boundary. For the saddle-point (2.16), we take the following ansatz for the vector field
@@
which is a second-order differential equation of $f_{y}(y)$ and is solved by the following general solution
@@
with integer valued $n$. These vector fields are large diffeomorphisms since they have nontrivial behaviors at the asymptotic boundary $y\to+\infty$,
@@
One way to regularize the one-loop partition function is to deform the extremal background by turning on slightly the temperature of the black hole, which in turn will corrects slightly the eigenvalues of the zero modes. In the next subsection, we will use the near extremal configuration given by (2.16) and (2.17) to calculate the leading corrections to the eigenvalues of the zero modes. We will see that these slightly corrected nonvanishing eigenvalues alter the partition function dramatically with a leading low-temperature effect.
@@ 2.3 Quantum corrections to the entropy
Note that $\delta\Lambda_{n}$ are small corrections to the eigenvalues which depend linearly on $T$, for nonzero modes with $\bar{\Lambda}_{n}\neq 0$, the corrections are purely polynomial in $T$,
@@
and correspondingly produce $\mathcal{O}(T)$ correction to the black hole mass at low temperature. So zero modes are the essential objects to resolve the mass gap puzzle addressed in Sec. 2.1.
@@
where we used the fact that $\bar{\Delta}[\bar{g}]$ is self-adjoint, i.e., $\langle h^{(m)}|\bar{\Delta}[\bar{g}]\delta h^{(n)}\rangle=\langle\bar{\Delta}[\bar{g}]h^{(m)}|\delta h^{(n)}\rangle$ and $h^{(n)}$ are orthonormalized. It is worthwhile to mention that the conical singularities induced by the improper range of the polar angle $\phi$ are inherited to the near horizon extremal geometry (2.16) through (2.15). However this effect will not affect the calculation of the eigenvalue corrections (2.45). The range of the polar angle $\varphi$ is canceled given that the zero modes satisfying the orthonormal condition (2.38). The integration in (2.45) is highly nontrivial. One way to carry out the integrand is to substitute the full perturbed metric $\bar{g}+\delta g$ into the linearized kinetic operator (2.27) and keep the leading $\mathcal{O}(T)$ term. The variation of the operator itself is intractable but it simplifies dramatically when contracting with the zero modes. Here we print the result we find
@@
$T_{q}$ can be viewed as the critical temperature below which the low-temperature quantum corrections become important. From (2.49), it is clear that the critical temperature rises as the acceleration of the black hole increases. However, this is not necessarily indicating that the high temperature quantum corrections are significant for a black hole with large acceleration. The mechanism presented here is working for the first-order low-temperature perturbation theory. The corrections from the levels that beyond the first order are still unknown at present stage.
@@ 3 Accelerating Reissner-Nordström black hole
$m$, $q$, $\alpha$ are the black hole mass, electric charge and acceleration parameters, respectively. $b$ is the time scaling parameter introduced for achieving consistent black hole thermodynamics Anabalon:2018qfv . $\Phi$ is the electric potential at the horizon, which ensures that the electric gauge field $A$ vanishes on the horizon. The accelerating of the charged black hole is induced by the cosmic string attached on the north ($\theta=0$) and south ($\theta=\pi$) poles. The string tensions are represented by the conical defects at each poles. Following the same argument in Sec. 2, the string tensions at the north $\mu_{+}$ and south $\mu_{-}$ poles can be expressed as
@@ 3.1 Near extremal thermodynamics and geometry
The near extremal thermodynamics fucus on the situation that the two horizon radius $r_{+}$ and $r_{-}$ depart a little bit from their extremal value $r_{0}$. Similar to the rotating case, there are three independent thermodynamic quantities which can be chosen as the temperature $T$, electric charge $Q$ and net string tension $\mu$. In the near extremal expansions, we take the temperature as a small parameter and keep the electric charge $Q_{0}=r_{0}$ and the net string tension $0<\mu_{0}=r_{0}\alpha<1$ fixed at their extremal values.
@@
The quantum corrections near extremality are introduced by using the Euclidean path integral with the saddle-point chosen as the extremal configuration. To find the near horizon near extremal configuration, we therefore perform the following coordinate transformation
@@
The metric (3.15) characterize the near horizon extremal geometry of the accelerating Reissner-Nordström black hole, which is a direct product of AdS2 and S2 with warping factors. Equation (3.16) is the extremal gauge field, which vanishes explicitly on the horizon $y=1$. When combined with the subleading terms (3.17) and (3.18), we get the near horizon near extremal configurations $d\bar{s}^{2}+\delta(ds^{2})$ and $\bar{A}+\delta(A)$ up to $\mathcal{O}(T)$.
@@ 3.2 Euclidean path integral near extremality and zero modes
which necessitates the joint diagonalization of the operators $\bar{\Delta}[\bar{g},\bar{A}]$, $\bar{P}[\bar{g},\bar{A}]$, and $\bar{O}_{\mathrm{int}}[\bar{g},\bar{A}]$. Inspired by the previous accelerating Kerr black hole case, we will focus on the path integral over zero modes of the saddle-points (3.15) and (3.16), which hopefully produces the leading quantum correction to the partition function at low temperature. Terminologically, the zero modes of the matrix operator in (3.32) can be divided into graviton zero modes and photon zero modes.
@@
with an arbitrary constant $C_{1}$, that is regular on the horizon and vanishes for $n=0$. Note that in deriving the above solution, we do not resort to the form of the function $f_{\phi}(y)$, which is different form the rotating black hole case analyzed in Sec. 2.2. This is highly due to the near horizon extremal geometry (3.15) is simply a direct product of AdS2 and S2 with warping factors. The function $f_{\phi}(y)$ can be solved independently from
@@
with integer valued $n$. These vector fields are indeed large diffeomorphisms since they do not die off at the asymptotic boundary $y\to\infty$ and thus can not be gauged away. The existence of the two arbitrary constants $C_{1}$ and $C_{2}$ indicates that there are two linearly independent graviton zero modes.
@@
The $n=0,\pm 1$ modes are excluded because the corresponding vector fields are Killing vectors on the background and do not produce fluctuations. The graviton zero modes which are generated by the $\xi$ with $C_{1}=0$ and $C_{2}\neq 0$ are called vector modes, whose explicit forms can be written as
@@
Note that for the vector modes, we have $|n|\geq 1$. The $n=0$ mode vanishes due to the fact that $\mathop{}\!\partial_{\phi}$ is the rotational Killing vector of the background. One can check that both (3.40) and (3.42) belong to the kernel of the operator $\bar{\Delta}[\bar{g},\bar{A}]$. It is worth noting that in the absence of acceleration, the Reissner-Nordström black hole has three sets of vector modes since there are three independent Killing vectors on S2 Banerjee:2023quv ; Blacker:2025zca . The presence of acceleration deforms the S2 and kills two of the Killing vectors, leaving only the translational symmetry along $\phi$. So there is only one set of the vector modes (3.42) in the accelerating case.
@@
It is straightforward to check that (3.47) is in the kernel of the operator $\bar{P}[\bar{g},\bar{A}]$. For a torsion free theory, the photon zero modes determined by (3.44) always give symmetric $\bar{\nabla}_{\mu}a_{\nu}$. It is obvious that the mixing term (3.31) vanishes for such zero modes. So the previous derived tensor, vector and photon zero modes indeed belong to the kernel of the matrix operator in (3.32).
@@ 3.3 Quantum corrections to the entropy
This will perturb the operators $\bar{\Delta}$ and $\bar{P}-(\bar{O}_{\mathrm{int}}^{\dagger}\bar{\Delta}^{-1}\bar{O}_{\mathrm{int}})/4$ and correct the the eigenvalues of the zero modes to nonzero ones. According to the perturbation theory at first order, the eigenvalue correction can be evaluated as the expectation value of the perturbed operator with respect to the original orthonormal eigenstates. However, from (3.31), it is obvious that the photon zero modes (3.47) satisfy
@@
So the variation of the operator $(\bar{O}_{\mathrm{int}}^{\dagger}\bar{\Delta}^{-1}\bar{O}_{\mathrm{int}})/4$ will not contribute to the first-order eigenvalue correction. Thus, to figure out the eigenvalue corrections, we only need to consider the variations of the operators $\bar{\Delta}$ and $\bar{P}$. Given the orthonormalized zero modes (3.40), (3.42) and (3.47), the corrections to their eigenvalues can be formulated as
@@
$T_{q}^{t}=\frac{r_{0}(3+\mu_{0}^{2})}{(1-\mu_{0}^{2})\sqrt{1+\mu_{0}^{2}}}\,,~~T_{q}^{v}=\frac{r_{0}(1-\mu_{0}^{2})(\mathrm{arctanh}\mu_{0}-\mu_{0})}{\mu_{0}^{3}\sqrt{1+\mu_{0}^{2}}}\,,~~T_{q}^{p}=\frac{r_{0}\mu_{0}}{\left(\mu_{0}-(1-\mu_{0}^{2})\mathrm{arctanh}\mu_{0}\right)\sqrt{1+\mu_{0}^{2}}}\,.$ (3.63)
@@
where $S_{0}$ is the extremal entropy. Note that the temperature dependence in the leading term of (3.64) is different from that in the accelerating Kerr case (2.50). This is because the absence of rotation and the presence of the gauge field each leads to additional zero modes and contributes $T^{1/2}$ factor in the partition function. Accordingly, the black hole thermal energy above extremality in this case is dominated by a term scaling as $(5/2)T$ at low temperature which replace it on the same footing as the Hawking radiation.
@@ 4 Accelerating Kerr-Newmann black hole
We finally analyze the low-temperature quantum corrections to the thermodynamics of the C-metric with both rotation and charge. As we will see, by turning on the rotation, the gravitational zero modes of vector type for the accelerating Reissner-Nordström black hole are no longer the case. The spacetime line element and the $U(1)$ gauge field for the accelerating Kerr-Newmann black hole in the Boyer-Lindquist coordinate system take the following forms Hong:2004dm ; Griffiths:2005se ; Griffiths:2005qp
@@
$m$, $a$, $q$, and $\alpha$ are the black hole mass, angular momentum, charge and acceleration parameters, respectively. $b$ is the time scaling parameter and $\Phi$ is the charge potential at the horizon. The outer and inner horizon radius $r_{+}$ and $r_{-}$ are now given by
@@ 4.1 Near extremal thermodynamics and geometry
As a consistency check, the chargeless limit $q_{0}\to 0$ of the above expansions recover the accelerating Kerr black hole case (2.10), (2.11), (2.12), and (2.13) while the static limit $q_{0}\to r_{0}$ reproduce the accelerating Reissner-Nordström black hole case (3.10), (3.11), (3.12), and (3.13).
To compute the low-temperature quantum corrections to the thermodynamic quantities, we need to figure out the near extremal configuration which can be obtained by the following coordinate transformation
@@
When substituting the transformations (4.14), (4.15) and (4.16) into the metric (4.1) and the gauge field (4.2) and taking the small-temperature limit with fixed angular momentum, charge and net string tension at their extremal values, we find the leading $\mathcal{O}(T^{0})$ terms